Please I want clear workings and formula used to arrive at the answer and detailed explanation for better understanding

A recent project nominated for consideration at your company has a four-year cash flow of $20,000; $25,000; $30,000; and $50,000. The cost of the project is $75,000.

a. If the required rate of return is 20%, conduct a discounted cash flow calculation to determine the NPV.

b. What is the benefit-cost ratio for the project?

c. What would the NPV of the above project be if the inflation rate was expected to be 4% in each of the next four years?
Answer #1

Part A

Investment

Year 1

Year 2

Year 3

Year 4

Cash flow

(75000)

20000

25000

30000

50000

PV@ 20%

1.00000

0.83333

0.69444

0.57870

0.48225

Present value

-75000

16667

17361

17361

24113

NPV

$502

Part B

Benefit-cost ratio = PV of benefit expected from the project / PV of the cost of the project = 75502/75000 = 1.01

Part C

Real discount rate = ((1.20/1.04)-1) = 15.38%

Investment

Year 1

Year 2

Year 3

Year 4

Cash flow

(75000)

20000

25000

30000

50000

PV@ 15.38%

1.00000

0.86670

0.75117

0.65104

0.56426

Present value

-75000

17334

18779

19531

28213

NPV

$8857